# Look Before You Leap: A GUI-Critic-R1 Model for Pre-Operative Error Diagnosis in GUI Automation

### v1

@@ Abstract
In recent years, Multimodal Large Language Models (MLLMs) have been extensively utilized for multimodal reasoning tasks, including Graphical User Interface (GUI) automation. Unlike general offline multimodal tasks, GUI automation is executed in online interactive environments, necessitating step-by-step decision-making based on the real-time status of the environment. This task has a lower tolerance for decision-making errors at each step, as any mistakes may cumulatively disrupt the process and potentially lead to irreversible outcomes like deletions or payments. To address these issues, we introduce a pre-operative critic mechanism that provides effective feedback prior to the actual execution, by reasoning about the potential outcome and correctness of actions. Specifically, we propose a Suggestion-aware Gradient Relative Policy Optimization (S-GRPO) strategy to construct our pre-operative critic model GUI-Critic-R1, incorporating a novel suggestion reward to enhance the reliability of the model’s feedback. Furthermore, we develop a reasoning-bootstrapping based data collection pipeline to create a GUI-Critic-Train and a GUI-Critic-Test, filling existing gaps in GUI critic data. Static experiments on the GUI-Critic-Test across both mobile and web domains reveal that our GUI-Critic-R1 offers significant advantages in critic accuracy compared to current MLLMs. Dynamic evaluation on GUI automation benchmark further highlights the effectiveness and superiority of our model, as evidenced by improved success rates and operational efficiency. The code is available at https://github.com/X-PLUG/MobileAgent/tree/main/GUI-Critic-R1.
@@ 1 Introduction
Recently, Multi-modal Large Language Models (MLLMs), leveraging their remarkable perception and reasoning capabilities, have demonstrated outstanding capabilities in various domains bai2025qwen2 ; wu2024deepseekvl . Among these, GUI automation, as a practical multi-modal application scenario, is emerging as a significant technological revolution in artificial intelligence interactions wang2024mobile2 ; wang2025mobile ; liu2025pc ; zhang2023appagent ; li2024appagent ; xu2024aguvis ; qin2025ui ; rawles2024androidworld . To be specific, given an online GUI device and a natural language instruction, it requires the GUI agent driven by MLLMs to generate a series of precise operations similar to how humans do zhang2024large ; tang2025survey , such as click, type, and scroll.
@@
Considering the above issues, in this paper, we propose a pre-operative critic (pre-critic) mechanism for GUI automation. Specifically, before performing an operation in online environments, the pre-critic mechanism first evaluates whether the operation generated by the agent is beneficial to the instruction completion, through comprehending the screenshot and analyzing potential results of the operation. Then, it provides real-time critical feedback to the agent, including the underlying causes of errors and corrective suggestions, to assist in refining its decision-making process. For example, as shown in Figure 1(b), given the instruction Rename the current audio to ‘my_audio’, the agent initially predicts the action of click (delete button) in the 5-th step. Such operation may cause the audio file to be deleted and the task to fail. Fortunately, with the pre-critic mechanism, we can catch the dangerous error before executing, and provide objective feedback to the agent. Besides, Figure 1(c) illustrates a case where the pre-critic identifies a more efficient method to turn on Bluetooth (i.e., enabling Bluetooth via the control center), thereby reducing the operation steps. We also exhibit some statistic results in Figure 1(d), where the pre-critic helps increase the success rate of baseline from 22.4% to 27.6% in the dynamic GUI automation rawles2024androidworld   [^1] More detailed experimental results about efficiency can be found in Table 2. In conclusion, these figures demonstrate that the introduction of pre-critic can effectively alleviate the aforementioned issues of error revision and operational inefficiency in online environments.
@@ 2 Related Work
LLM-based GUI Agent. Recently, significant attention has been directed toward Graphical User Interface (GUI) agents for task automation on smart devices designed for the mobile and web environments zhang2023appagent ; li2024appagent ; wang2024mobile2 ; zhang2024ufo ; zheng2025vem ; yan2023gpt ; hong2024cogagent ; zhang2024smartagent ; ma2024coco ; song2024visiontasker ; wen2024autodroid ; wen2023droidbot . Despite these advancements, GUI agents frequently make erroneous decisions. Some research has incorporated reflection modules in GUI automation frameworks to verify operation correctness based on the both current screenshot and the screenshot after execution wang2024mobile2 ; liu2025pc ; agashe2025agent . For example, Mobile-Agent-v2 frameworks wang2024mobile2 employ multi-agent architectures that separate planning, decision-making, and reflection to optimize task tracking, memory, and error correction. But these methods require additional steps to undo actions and pose the risk of irreversible operations, resulting in lower efficiency and accuracy. In this paper, we introduces a pre-critic mechanism for diagnosing potential errors.
@@ 3.1 Problem Definition of Pre-Operative Critic
In this paper, we propose a pre-operative critic model $\pi_{critic}(\epsilon,a)$ to critique the decision-making process of $\pi_{agent}$, before the action is actually executed. Specifically, taking state $\epsilon$ and action $a$ as inputs, the model produces a correctness score ${l}\in[0,1]$ that reflects the rightness of $a$. Besides, it generates critique ${c}$ and corrective suggestion ${s}$ in natural language, with the former explaining the rationale behind the correctness judgment and the latter suggesting a better action if $a$ is incorrect.
@@ 3.2 Data Collection Pipeline with Reasoning Bootstrapping
In this section, we detail the data collection pipeline of our proposed GUI-Critic-Train and GUI-Critic-Test datasets. Each sample in the dataset includes the environmental state $\epsilon$ of the step, a operation candidate $a$, as well as annotations for correctness score $l$ and suggestion ${s}$. This section commences by introducing the collection of operations in Section 3.2.1, followed by the presentation of our reasoning bootstrapping method for generating critique $c$ in Section 3.2.2.
@@ 3.2.1 GUI Operations Collection
To begin with, we collect successful automation trajectories from publicly accessible datasets, encompassing correct step-level operations across various GUI scenarios. The action in a specific state $\epsilon$ is considered as the operation candidate whose correctness score is positive ($l=1$), and the corresponding action description is considered as the suggestion ${s}$.
@@ 3.2.2 GUI Critiques Generation
Reasoning Bootstrapping. Following the above format, we leverage current MLLMs to generate CoT reasoning progress. However, performing reverse annotation of CoT conditioned on ground-truth score $l$ and suggestion $s$ is harmful. This is because that the MLLMs may overly depend on pre-known annotations, potentially biasing the thought process toward these outcomes rather than reflecting the actual critique reasoning sequence. Therefore, we propose a reasoning bootstrapping strategy to generate high-quality thoughts without prior knowledge of the ground-truth $l$ and $s$. Specifically, merely provided with environmental states $\epsilon$ and action candidates $a$, it forces the model to simulate a genuine critic reasoning process for ${max}$ times as follows:
@@ 3.3.1 RFT Cold-Start
$\small\mathcal{L}_{rft}=\mathbb{E}_{(\epsilon,a,l,s)\sim\mathcal{D}_{c\_action},\ (\hat{\epsilon},\hat{a},\hat{l},\hat{s},\hat{t})\sim\mathcal{D}_{c\_cot}}-\text{log}(\pi_{critic}(l,s|\epsilon,a))-\text{log}(\pi_{critic}(\hat{l},\hat{s},\hat{t}|\hat{\epsilon},\hat{a})).$ (2)
@@ 3.3.2 Suggestion-aware Group Relative Policy Optimization
where ${o}=({l},{s},{c})$ is the output of pre-critic model $\pi_{critic}$. $\mathbb{I}_{similar}(\cdot)$ calculates the similarity between the inputs leveraging large language models, and output a binary judgment. Based on Eq. (3), $r_{s}$ provides a quantitative and objective metric to assess the correctness of suggestions, thereby facilitating more precise evaluations of $\pi_{critic}$ efficacy. Moreover, this direct feedback enables models to learn from their outputs iteratively, refining their progressive CoT reasoning process to produce more accurate action critique and corrective suggestions.
@@
[…]
@@ 4.2.2 Dynamic Evaluation
Critic Accuracy Suggestion Accuracy Critic Accuracy Suggestion Accuracy
@@
$r_{f}$ $r_{a}$ $r_{s}$ Critic Accuracy Suggestion Accuracy Critic Accuracy Suggestion Accuracy
@@ B More Dynamic Evaluation
In this paper, we conduct the dynamic evaluation on AndroidWorld rawles2024androidworld and illustrate the experiment results in the main text (Section 4.2.2). In this section, we introduce more details of the implementation of the experiments and more experimental results.
@@ B.1 Supplementary Implementation Detials
We adopt the M3A (Multimodal Autonomous Agent for Android) framework introduced in AndroidWorld rawles2024androidworld as the backbone. In the M3A framework, the decision agent is provided with a list of available action types, guidelines for operating the phone, and a list of UI elements derived from the leaf nodes of the Android accessibility tree. The agent receives the current screenshot along with a Set-of-Mark annotated screenshot, which includes bounding boxes with numeric labels at the top-left corner of each UI element. At this stage, the agent attempts to execute the generated action, referencing specific marks. Prior to action execution, we insert a pre-operative critic to evaluate whether the agent’s proposed action is conducive to achieving the instruction. If the action is deemed correct, it proceeds to execution as normal. Conversely, if the action is considered incorrect, we input the critic model’s analytical recommendations to the decision agent, prompting it to reassess and formulate a new decision. After action execution, another agent is adopted to deliver a concise summary following the execution of the action. We leverage such semantic summaries to serve as a record of the action history. We follow the Androidworld to configure the UI element detection procedure and the definition of the action space.
@@ B.2 Supplementary Experiments
Table 5: Supplementary dynamic evaluation results on the AndroidWorld rawles2024androidworld benchmark with GPT-4o gpt4o as the baseline.
@@
Qwen2.5-VL-7B bai2025qwen2 16.9 32.4
Qwen2.5-VL-72B bai2025qwen2 25.9 36.8
GPT-4o gpt4o 24.0 43.1
@@
We extend the dynamic evaluation experiment to incorporate GPT-4o gpt4o as the decision agent. As evidenced in Table 5, employing GPT-4o gpt4o as the decision-making agent yields performance benefits over Qwen2.5-VL-72B bai2025qwen2 , underscoring GPT-4o’s robust perception and reasoning capabilities. The pre-operative critic improves accuracy and exemplifies the efficacy of the baseline in detecting erroneous decisions and offering constructive improvement suggestions. In comparison with other pre-operative models, our GUI-Critic-R1 exhibits superior performance outcomes. When utilizing GPT-4o as the baseline, the advantage of the pre-critic in terms of the Efficient Advantage Rate (EAR) becomes more pronounced. GPT-4o tends to engage in excessive exploration during GUI automation, often opting to attempt further actions rather than terminating early, thereby resulting in a greater number of steps. In contrast, the pre-critic effectively anticipates erroneous or redundant exploration, thereby reducing unnecessary steps and manifesting a superior EAR metric.

### v2

@@ Abstract
In recent years, Multimodal Large Language Models (MLLMs) have been extensively utilized for multimodal reasoning tasks, including Graphical User Interface (GUI) automation. Unlike general offline multimodal tasks, GUI automation is executed in online interactive environments, necessitating step-by-step decision-making based on the real-time status of the environment. This task has a lower tolerance for decision-making errors at each step, as any mistakes may cumulatively disrupt the process and potentially lead to irreversible outcomes like deletions or payments. To address these issues, we introduce a pre-operative critic mechanism that provides effective feedback prior to the actual execution, by reasoning about the potential outcome and correctness of actions. Specifically, we propose a Suggestion-aware Group Relative Policy Optimization (S-GRPO) strategy to construct our pre-operative critic model GUI-Critic-R1, incorporating a novel suggestion reward to enhance the reliability of the model’s feedback. Furthermore, we develop a reasoning-bootstrapping based data collection pipeline to create a GUI-Critic-Train and a GUI-Critic-Test, filling existing gaps in GUI critic data. Static experiments on the GUI-Critic-Test across both mobile and web domains reveal that our GUI-Critic-R1 offers significant advantages in critic accuracy compared to current MLLMs. Dynamic evaluation on GUI automation benchmark further highlights the effectiveness and superiority of our model, as evidenced by improved success rates and operational efficiency. The code is available at https://github.com/X-PLUG/MobileAgent/tree/main/GUI-Critic-R1.
@@ 1 Introduction
Recently, Multi-modal Large Language Models (MLLMs), leveraging their remarkable perception and reasoning capabilities, have demonstrated outstanding capabilities in various domains bai2025qwen2; wu2024deepseekvl. Among these, GUI automation, as a practical multi-modal application scenario, is emerging as a significant technological revolution in artificial intelligence interactions wang2025mobile; liu2025pc; zhang2023appagent; li2024appagent; xu2024aguvis; qin2025ui; rawles2024androidworld. To be specific, given an online GUI device and a natural language instruction, it requires the GUI agent driven by MLLMs to generate a series of precise operations similar to how humans do zhang2024large; tang2025survey, such as click, type, and scroll.
@@
Considering the above issues, in this paper, we propose a pre-operative critic (pre-critic) mechanism for GUI automation. Specifically, before performing an operation in online environments, the pre-critic mechanism first evaluates whether the operation generated by the agent is beneficial to the instruction completion, through comprehending the screenshot and analyzing potential results of the operation. Then, it provides real-time critical feedback to the agent, including the underlying causes of errors and corrective suggestions, to assist in refining its decision-making process. For example, as shown in Figure 1(b), given the instruction Rename the current audio to ‘my_audio’, the agent initially predicts the action of click (delete button) in the 5-th step. Such operation may cause the audio file to be deleted and the task to fail. Fortunately, with the pre-critic mechanism, we can catch the dangerous error before executing, and provide objective feedback to the agent. Besides, Figure 1(c) illustrates a case where the pre-critic identifies a more efficient method to turn on Bluetooth (i.e., enabling Bluetooth via the control center), thereby reducing the operation steps. We also exhibit some statistic results in Figure 1(d), where the pre-critic helps increase the success rate of baseline from 22.4% to 27.6% in the dynamic GUI automation rawles2024androidworld ***More detailed experimental results about efficiency can be found in Table 2. In conclusion, these figures demonstrate that the introduction of pre-critic can effectively alleviate the aforementioned issues of error revision and operational inefficiency in online environments.
@@ 2 Related Work
LLM-based GUI Agent. Recently, significant attention has been directed toward Graphical User Interface (GUI) agents for task automation on smart devices designed for the mobile and web environments zhang2023appagent; li2024appagent; wang2024mobile; zhang2024ufo; zheng2025vem; yan2023gpt; hong2024cogagent; zhang2024smartagent; ma2024coco; song2024visiontasker; wen2024autodroid; wen2023droidbot. Despite these advancements, GUI agents frequently make erroneous decisions. Some research has incorporated reflection modules in GUI automation frameworks to verify operation correctness based on the both current screenshot and the screenshot after execution wang2024mobile2; liu2025pc; agashe2025agent. For example, Mobile-Agent-v2 frameworks wang2024mobile employ multi-agent architectures that separate planning, decision-making, and reflection to optimize task tracking, memory, and error correction. But these methods require additional steps to undo actions and pose the risk of irreversible operations, resulting in lower efficiency and accuracy. In this paper, we introduces a pre-critic mechanism for diagnosing potential errors.
@@ 3.1 Problem Definition of Pre-Operative Critic
In this paper, we propose a pre-operative critic model $\pi_{critic}(\epsilon,a)$ to critique the decision-making process of $\pi_{agent}$, before the action is actually executed. Specifically, taking state $\epsilon$ and action $a$ as inputs, the model produces a correctness score ${l}\in[0,1]$ that reflects the rightness of $a$. Besides, it generates critique c{c} and corrective suggestion s{s} in natural language, with the former explaining the rationale behind the correctness judgment and the latter suggesting a better action if $a$ is incorrect.
@@ 3.2 Data Collection Pipeline with Reasoning Bootstrapping
In this section, we detail the data collection pipeline of our proposed GUI-Critic-Train and GUI-Critic-Test datasets. Each sample in the dataset includes the environmental state $\epsilon$ of the step, a operation candidate $a$, as well as annotations for correctness score $l$ and suggestion s{s}. This section commences by introducing the collection of operations in Section 3.2.1, followed by the presentation of our reasoning bootstrapping method for generating critique $c$ in Section 3.2.2.
@@ 3.2.1 GUI Operations Collection
To begin with, we collect successful automation trajectories from publicly accessible datasets, encompassing correct step-level operations across various GUI scenarios. The action in a specific state $\epsilon$ is considered as the operation candidate whose correctness score is positive ($l=1$), and the corresponding action description is considered as the suggestion s{s}.
@@ 3.2.2 GUI Critiques Generation
Reasoning Bootstrapping. Following the above format, we leverage current MLLMs to generate CoT reasoning progress. However, performing reverse annotation of CoT conditioned on ground-truth score $l$ and suggestion $s$ is harmful. This is because that the MLLMs may overly depend on pre-known annotations, potentially biasing the thought process toward these outcomes rather than reflecting the actual critique reasoning sequence. Therefore, we propose a reasoning bootstrapping strategy to generate high-quality thoughts without prior knowledge of the ground-truth $l$ and $s$. Specifically, merely provided with environmental states $\epsilon$ and action candidates $a$, it forces the model to simulate a genuine critic reasoning process for max{max} times as follows:
@@ 3.3.1 RFT Cold-Start
$).\small\mathcal{L}_{rft}=\mathbb{E}_{(\epsilon,a,l,s)\sim\mathcal{D}_{c\_action},\ (\hat{\epsilon},\hat{a},\hat{l},\hat{s},\hat{t})\sim\mathcal{D}_{c\_cot}}-\text{log}(\pi_{critic}(l,s|\epsilon,a))-\text{log}(\pi_{critic}(\hat{l},\hat{s},\hat{t}|\hat{\epsilon},\hat{a}$)). (2)
@@ 3.3.2 Suggestion-aware Group Relative Policy Optimization
where o=(l,s,c){o}=({l},{s},{c}) is the output of pre-critic model $\pi_{critic}$. $\mathbb{I}_{similar}(\cdot)$ calculates the similarity between the inputs leveraging large language models, and output a binary judgment. Based on Eq. (3), $r_{s}$ provides a quantitative and objective metric to assess the correctness of suggestions, thereby facilitating more precise evaluations of $\pi_{critic}$ efficacy. Moreover, this direct feedback enables models to learn from their outputs iteratively, refining their progressive CoT reasoning process to produce more accurate action critique and corrective suggestions.
@@
$\small\mathcal{L}_{\mathrm{GRPO}}=\frac{1}{|G|}\sum_{i=1}^{G}\min\!\Bigl(\frac{\pi_{critic}(o_{i}\mid\epsilon,a)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid\epsilon,a)}\,A_{i},\,\\ \mathrm{clip}\!\Bigl(\frac{\pi_{critic}(o_{i}\mid\epsilon,a)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid\epsilon,a)},1-\varepsilon,\,1+\varepsilon\Bigr)A_{i}\Bigr)-\,\beta\,D\bigl(\pi_{critic}\,\big\|\,\pi_{\mathrm{ref}}\bigr),$ (4)
@@ 4.2.2 Dynamic Evaluation
Critic
Accuracy
Suggestion
Accuracy
Critic
Accuracy
Suggestion
Accuracy
@@
$r_{f}$ $r_{a}$ $r_{s}$
Critic
Accuracy
Suggestion
Accuracy
Critic
Accuracy
Suggestion
Accuracy
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grants 62322212, U23A20387), the Beijing Natural Science Foundation (No. L221013), and the CAS Project for Young Scientists in Basic Research (YSBR-116).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s contributions and scope.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The limitations are discussed in the Appendix.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. 4.
    Experimental Result Reproducibility
  13. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  14. Answer: [Yes]
  15. Justification: We have provided the necessary information for reproduction in Section LABEL:sec:experiment-setup.
  16. 5.
    Open access to data and code
  17. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  18. Answer: [No]
  19. Justification: Justification: The codes for reproducing all experimental results will be released as soon as the paper is accepted by NeurIPS.
  20. 6.
    Experimental Setting/Details
  21. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  22. Answer: [Yes]
  23. Justification: We have provided the training and test details necessary to understand the results in Section 4.1.
  24. 7.
    Experiment Statistical Significance
  25. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  26. Answer: [Yes]
  27. 8.
    Experiments Compute Resources
  28. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  29. Answer: [Yes]
  30. Justification: We provide the compute workers details in Section 4.1, and the amount of compute are illustrated in Appendix.
  31. 9.
    Code Of Ethics
  32. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  33. Answer: [Yes]
  34. Justification: The research conforms with the NeurIPS Code of Ethics in every respect.
  35. 10.
    Broader Impacts
  36. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  37. Answer: [N/A]
  38. Justification: The paper prioritizes discussing technical methodologies and results rather than societal impacts.
  39. 11.
    Safeguards
  40. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  41. Answer: [N/A]
  42. Justification: The paper exploring GUI automation on public datasets poses no such risks.
  43. 12.
    Licenses for existing assets
  44. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  45. Answer: [Yes]
  46. Justification: We have properly credited the used assets. All the datasets introduced in this paper are available for free to researchers for non-commercial use.
  47. 13.
    New Assets
  48. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  49. Answer: [N/A]
  50. Justification: The codes will be released once upon the paper’s acceptance.
  51. 14.
    Crowdsourcing and Research with Human Subjects
  52. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  55. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  56. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  57. Answer: [N/A]
  58. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  59. 16.
    Declaration of LLM usage
  60. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  61. Answer: [Yes] .
  62. Justification: The paper uses LLM and describes in detail the type of model used (Sec 4.1) and prompt in Appendix.
@@ B More Dynamic Evaluation
In this paper, we conduct the dynamic evaluation on AndroidWorld [rawles2024androidworld] and illustrate the experiment results in the main text (Section 4.2.2). In this section, we introduce more details of the implementation of the experiments and more experimental results.
@@ B.1 Supplementary Implementation Detials
We adopt the M3A (Multimodal Autonomous Agent for Android) framework introduced in AndroidWorld [rawles2024androidworld] as the backbone. In the M3A framework, the decision agent is provided with a list of available action types, guidelines for operating the phone, and a list of UI elements derived from the leaf nodes of the Android accessibility tree. The agent receives the current screenshot along with a Set-of-Mark annotated screenshot, which includes bounding boxes with numeric labels at the top-left corner of each UI element. At this stage, the agent attempts to execute the generated action, referencing specific marks. Prior to action execution, we insert a pre-operative critic to evaluate whether the agent’s proposed action is conducive to achieving the instruction. If the action is deemed correct, it proceeds to execution as normal. Conversely, if the action is considered incorrect, we input the critic model’s analytical recommendations to the decision agent, prompting it to reassess and formulate a new decision. After action execution, another agent is adopted to deliver a concise summary following the execution of the action. We leverage such semantic summaries to serve as a record of the action history. We follow the Androidworld to configure the UI element detection procedure and the definition of the action space.
@@ B.2 Supplementary Experiments
Table 5: Supplementary dynamic evaluation results on the AndroidWorld [rawles2024androidworld] benchmark with GPT-4o [gpt4o] as the baseline.
@@
Qwen2.5-VL-7B [bai2025qwen2] 16.9 32.4
Qwen2.5-VL-72B [bai2025qwen2] 25.9 36.8
GPT-4o [gpt4o] 24.0 43.1
@@
We extend the dynamic evaluation experiment to incorporate GPT-4o [gpt4o] as the decision agent. As evidenced in Table 5, employing GPT-4o [gpt4o] as the decision-making agent yields performance benefits over Qwen2.5-VL-72B [bai2025qwen2], underscoring GPT-4o’s robust perception and reasoning capabilities. The pre-operative critic improves accuracy and exemplifies the efficacy of the baseline in detecting erroneous decisions and offering constructive improvement suggestions. In comparison with other pre-operative models, our GUI-Critic-R1 exhibits superior performance outcomes. When utilizing GPT-4o as the baseline, the advantage of the pre-critic in terms of the Efficient Advantage Rate (EAR) becomes more pronounced. GPT-4o tends to engage in excessive exploration during GUI automation, often opting to attempt further actions rather than terminating early, thereby resulting in a greater number of steps. In contrast, the pre-critic effectively anticipates erroneous or redundant exploration, thereby reducing unnecessary steps and manifesting a superior EAR metric.
B.3 Additional Case Study